OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3 (x )

OpenStudy (johnweldon1993):

remember the equation of a circle \[(x - h)^2 + (y - k)^2 = r^2\] where (h,k) is your center and 'r' is your radius

OpenStudy (anonymous):

\[x ^{2}+y ^{2}-10x-12y+52=0\] ans